Shift report 20020109 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Jan 09 2002 - 16:38:13 EST

  • Next message: Flemming Videbaek: "RHIC - scheduling physicist (bad news)"

    ------------------------------------------------------------
    Date: Wed Jan  9 16:38:13 2002, Shift: 08:00-16:00
    Supervisor: KH, with: EJK
    Log book#12 pages 33 - 33
    
    Shift Summary:
    ------------------------------------------------------------
    No beam during entire shift.
    There was an access to allow pp2pp to install their
    detectors.  We took advantage of that opportunity to
    do some things in the IR.
    
    1. The lead shield in front of BBL #43 was removed as it
       was no longer needed.
    2. The timing of the inelastic triggers was measured and
       found was found to be off by 8ns, exactly the 
       amount moved yesterday by adjusting with beam.  
       As it turned out the beam yesterday afternoon had a 
       very bad longitudinal emittance, so there were very 
       long bunches and we got fooled by what we saw on the
       scope and probably picked up some halo source that 
       was offset by that amount.  The result is that we 
       set the timing to collect on the wrong bucket and 
       it explains the strange vertex distributions 
       observed during the night.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Jan 09 2002 - 16:38:59 EST